iosangel:

The prime factoriztion of 600 is 2 ^a 3^b 5^c. what is the value of A?

Vocaloid:

Keep dividing 600 into factors until you get to primes only 600 = 5*120 120 = 5*24 So far that gives us 5*5 or 5^2 as one factor Using similar logic, break down 24 into factors of 2 and 3 only. It’s asking for the exponent on 2^a so you simply need to find out how many 2’s you can get in your factorization.